Eight times a number plus 6 less than twice the number is 34
Answer by integral13(8)

To solve this equation:
8 times a number plus 6 less than twice the number is 34.
8x+6-2x=34
Subtract 8x and 2x and you get 6x.
6x+6=34
Subtract 6 from BOTH SIDES, leaving you with 34-6=28
The equation...
6x=28
Divide six on BOTH SIDES
28/6= 14/3 OR 4.666666
x=4.666666 or 14/3
Hope that Helps!
Integral 13